A varies directly as h. A = 48 when h = 4. Find h when A = 36
Answer by ankor@dixie-net.com(12704)

A varies directly as h. A = 48 when h = 4. Find h when A = 36
:
A direct variation equation can take the form A = k*h
:
Substitute for A and h to find k
48 = 4k
k = 48/4
k = 12
:
Now our equation can be written:
A = 12h
:
Find h when A = 36
Substitute 36 for A and find h
36 = 12h
h = 36/12
h = 3